OpenStudy (anonymous):

Find the sum of the measures of the angles in a nonagon (9 sides). a. 620° b. 900° c. 1260° d. 1620°

OpenStudy (johnweldon1993):

We can us the equation \[(n-2) \times 180\] Where 'n' is the number of sides of the polygon...here we have 9 sides ...so \[(9-2) \times 180\] \[7 \times 180 = \space?\]

OpenStudy (anonymous):

c

OpenStudy (johnweldon1993):

There you go!